# Logical coherence in 2D compass codes

### v1

@@ Abstract
where the approximations are valid at large $l$, where $l_{h}\gg t_{h}$, thus $l_{h}\approx\frac{l}{h}$. In the regime where $\epsilon_{l}(h\theta),\lambda_{l}(h\theta)$ are both sufficiently small, we can apply similar approximations as above for the X-Shor code. The resulting measure of coherence is approximately $\kappa_{\mathcal{C}_{l,h}}\approx\frac{l}{h}\kappa_{l}(h\theta)$. This result tells us that the coherence threshold for $\mathcal{C}_{l,h}$ is $\frac{\pi}{2h}$ and that the exponential suppression of loss of fidelity and coherence is slowed down by a linear factor of $l/h$. Bounding the rotated surface code.— Under MWPM decoding, the coherence thresholds $\pi/4$ and $\pi/6$ of the two families, $\mathcal{C}_{l,2}$ and $\mathcal{C}_{l,3}$, respectively, serve as bounding cases of the $\pi/5$ coherence threshold of the surface code which we reproduced as shown in Fig. 2. However, based on the results of Venn, Behrends, and Béri [25] under ML decoding, the surface code might be able to reach the threshold of $(0.28\pm 0.01)\pi$, which is slightly above the threshold of $\mathcal{C}_{l,2}$. While an exact explanation is left for future work, the number of X and Z stabilizer generators being the same for the surface code and the $\mathcal{C}_{l,2}$ is suggestive of an underlying structural reason for the proximity of the two codes’ thresholds.
@@
FIGURE 2: Infidelity thresholds $\theta_{th}^{r_{1}}$ of different compass codes. Random compass codes (RCC) are interpolated between the Z-Shor code (ZSC) and the X-Shor code (XSC) parametrized by the X-check density $q_{shor}$ using numerical experiments to estimate the average threshold. Each data point for RCC uses 200 random codes and 200 samples for each distance across $d=9,13,17,21$. The threshold ranges are derived from plots in Appendix A. We display the analytically determined thresholds for the Z-stacked Shor code families $\mathcal{C}_{l,2}$ and $\mathcal{C}_{l,3}$, and the numerical values for the rotated surface code (RSC MWPM) based on our simulation, which agrees with Bravyi et al.’s result [11]. Finally, the ML decoding upper bound based on [25] is displayed as RSC ML.
@@
Finding a smoother transition might provide more insight into an analytical pathway toward a constant lower bound for the surface code threshold. We find that – at least on average – it is possible to smoothly interpolate between the extreme thresholds of the rotated surface code and the Z-Shor code thresholds using random compass codes of different X-check densities (Fig. 2). We simulate random compass code families parameterized with $q_{shor}$, where $q_{shor}=1.0$ corresponds to the X-Shor code and $q_{shor}=0.0$ to the Z-Shor code. The parameter represents the portion of the coloring that comprises X-type (blue) stabilizers. In this parameter manifold $\mathcal{C}_{l,3}$ is $q_{shor}=1/3$ and both $\mathcal{C}_{l,2}$ and the rotated surface code are at $q_{shor}=1/2$. For the numerical experiments, we implemented and generalized for compass codes a Majorana mode simulator invented by Bravyi et al. [11]. Our code, msim, is open sourced [40]. For MWPM decoding, we used PyMatching [41].
@@
Conclusions and outlook.— 2D compass codes are promising for exploring the open problem of proving a constant analytical lower bound for the rotated surface code threshold under uniform coherent Z-rotation, the worst-case scenario for coherent noise processes. We showed how the Z-Shor code can suppress coherent error in the $\sin(\theta)<1/L$ regime, similar to the toric code [7]. Given that the Z-Shor code does not have a threshold and numerical evidence by Bravyi et al. [11] points to a constant lower bound for the rotated surface code, there is room for improvement analytically as well. We introduced the Z-stacked Shor code families that afford approximate analytical analysis of the space. Two of these families yield thresholds near the rotated surface code value. Numerical investigations via the Majorana mode simulator can now be extended to compass codes with our open-source tool, msim [40]. These simulations show that it is possible to find structural properties that move the threshold on average. Still, the simple X-check density, $q_{shor}$, itself fails to have a high enough resolution of the threshold to fully explain the relationship between the two bounding cases of stacked Z-Shor codes and the rotated surface code.
@@
The authors thank D. Debroy and R. Calderbak for useful discussions. This work was supported by ARO/LPS QCISS program (W911NF-21-1-0005) and the NSF QLCI for Robust Quantum Simulation(OMA-2120757).
Code and data are available for download at [42].
@@ Appendix A Supporting data for threshold calculations
FIGURE 3: The underlying data for the threshold plot in the main text. We manually determine the upper and lower bounds for the thresholds by picking a regime that is clearly below threshold and one that is clearly above. The bounds are denoted with the two blue vertical lines as well as the numerical values in the highlighted box. Rows are for the same code families, while columns represent the single-use channel fidelity $r_{1}$ (left column) and the diamond distance $d_{\diamond}$ (right column).

### v2

@@ III.4 Coherence threshold for the $Z$-stacked Shor codes
where the approximations are valid at large $l$, where $l_{h}\gg t_{h}$, thus $l_{h}\approx\frac{l}{h}$. In the regime where $\epsilon_{l}(h\theta),\lambda_{l}(h\theta)$ are both sufficiently small, we can apply similar approximations as above for the $X$-Shor code. The resulting measure of coherence is approximately $\kappa_{\mathcal{C}_{l,h}}\approx\frac{l}{h}\kappa_{l}(h\theta)$. This result tells us that the coherence threshold for $\mathcal{C}_{l,h}$ is $\frac{\pi}{2h}$ and that the exponential suppression of infidelity and coherence is slowed down by a linear factor of $l/h$.
In the case of the $Z$-stacked Shor codes, not surprisingly, our calculations reflect performance under MWPM decoding (see Appendix B for numerical evidence).
III.5 Bounding the rotated surface code
Under MWPM decoding, the infidelity thresholds $\pi/4$ and $\pi/6$ of the two families, $\mathcal{C}_{l,2}$ and $\mathcal{C}_{l,3}$, respectively, serve as bounding cases of the $\pi/5$ infidelity threshold of the surface code under MWPM decoding, which we reproduced as shown in Fig. 4.
@@ III.5 Bounding the rotated surface code
FIGURE 3: Infidelity thresholds $\theta_{th}^{r_{1}}$ of different compass codes. Random compass codes (RCC) are interpolated between the $Z$-Shor code (ZSC) and the $X$-Shor code (XSC) parametrized by the $X$-check density $q_{shor}$ using numerical experiments to estimate the average threshold. We display the analytically determined thresholds for the $Z$-stacked Shor code families $\mathcal{C}_{l,2}$ and $\mathcal{C}_{l,3}$, and the numerical values for the rotated surface code (RSC MWPM) based on our simulation, which agrees with Bravyi et al.’s result [11]. Finally, the ML decoding upper bound based on [25] is displayed as RSC ML.
@@
Finding a smoother transition might provide more insight into an analytical pathway toward a constant lower bound for the surface code threshold. We find that – at least on average – it is possible to smoothly interpolate between the extreme thresholds of the rotated surface code and the $Z$-Shor code thresholds using random compass codes of different $X$-check densities (Fig. 3). We simulate random compass code families parameterized with $q_{shor}$, where $q_{shor}=1.0$ corresponds to the $X$-Shor code and $q_{shor}=0.0$ to the $Z$-Shor code. The parameter represents the portion of the coloring that comprises $X$-type stabilizers. In this parameter manifold $\mathcal{C}_{l,3}$ is $q_{shor}=1/3$ and both $\mathcal{C}_{l,2}$ and the rotated surface code are at $q_{shor}=1/2$. We evaluated the infidelity threshold for these codes numerically. Each data point uses 200 random codes and 200 samples for each distance across $d=9,13,17,21$. The threshold ranges are derived from plots in Appendix A. For these experiments, we implemented and generalized for compass codes a Majorana mode simulator invented by Bravyi et al. [11]. Our simulator, msim, is open sourced [41]. For MWPM decoding, we used PyMatching[42].
IV Conclusions and outlook
@@
2D compass codes are promising for exploring the open problem of proving a constant analytical lower bound for the rotated surface code threshold under uniform coherent $Z$-rotation, the worst-case scenario for coherent noise processes. We showed how the $Z$-Shor code can suppress coherent error in the $\sin(\theta)<1/L$ regime, similar to the toric code [7]. Given that the $Z$-Shor code does not have a threshold and numerical evidence by Bravyi et al. [11] points to a constant lower bound for the rotated surface code, there is room for improvement analytically as well.
We introduced the $Z$-stacked Shor code families that afford analytical analysis of the space. Two of these families yield thresholds under minimum weight decoding that provide upper and lower bounds to the rotated surface code’s infidelity threshold. While an exact explanation is left for future work, the number of X and Z stabilizer generators being the same for the surface code and the $\mathcal{C}_{l,2}$ is suggestive of an underlying structural reason for the proximity of the two codes’ thresholds.
Numerical investigations via the Majorana mode simulator can now be extended to compass codes with our open-source tool, msim [41]. These simulations show that it is possible to find structural properties that move the threshold on average. Still, the simple $X$-check density, $q_{shor}$, in itself, on average fails to have a high enough resolution of the threshold to fully explain the relationship between the two bounding cases of stacked $Z$-Shor codes and the rotated surface code.
Based on the results of Venn, Behrends, and Béri [25] and the author’s other work [40] under ML decoding, the surface code is able to reach the threshold of $(0.28\pm 0.01)\pi$. This is slightly above the threshold of $\mathcal{C}_{l,2}$ under MWPM decoding. We conjecture that an ML decoder will not be able to increase a higher threshold in the case of these codes, despite potentially achieving lower error rates. Thus, bounding the surface code under ML decoding using compass codes is left for future work.
@@
The authors thank D. Debroy and R. Calderbak for useful discussions. This work was supported by ARO/LPS QCISS program (W911NF-21-1-0005) and the NSF QLCI for Robust Quantum Simulation(OMA-2120757). Code and data are available for download at [43].
@@ Appendix A Supporting data for threshold calculations
In Fig. 4, we display the data behind our threshold calculations based on random code families and the rotated surface code. For the random code families, each datapoint consists of 200 random families with the same $X$ check density and 200 samples from each families. Given the logical infidelity $1-\cos(\theta_{s})$ for each measured syndrome with logical rotation $\theta_{s}$, we display the mean and the standard error for the observed $4\times 10^{4}$ logical infidelities. We compare the average logical infidelity with the average logical diamond distance $2|\sin(\theta_{s})|$ and standard error of the diamond distance. We observe slightly higher threshold values for the diamond distance.
FIGURE 4: The underlying data for the threshold plot in the main text. We manually determine the upper and lower bounds for the thresholds by picking a regime that is clearly below threshold and one that is clearly above based on intersections observed at distances 9,13,17,21. The bounds are displayed using two blue vertical lines as well as the numerical values in the highlighted box. Rows are for the same code families, while columns represent the single-use channel fidelity $r_{1}$ (left column) and the diamond distance $d_{\diamond}$ (right column).
Appendix B Numerical verification of analytical results
Here, we compare the predictions from our analytical results for repetition code based compass codes with numerical simulation data based on msim [41] using Pymatching [42] for MWPM decoding. There is perfect agreement between the two methods when the sample size is sufficiently large, verifying our calculations as well as our generalization of the Majorana simulator to the 2D compass codes.
FIGURE 5: Logical infidelity $r_{1}$ for repetition codes. Exact (min weight) line shows the calculation using the repetition code based formulas in the text, without any approximations. For the numerical data, each data point is based on $10^{4}$ samples.
FIGURE 6: Logical infidelity $r_{1}$ for $X$-Shor codes. Exact (min weight) line shows the calculation using the repetition code based formulas in the text, without any approximations. For the numerical data, each data point is based on $10^{4}$ samples.
FIGURE 7: Logical infidelity $r_{1}$ for $Z$-Shor codes in a high error regime, in an oscillatory regime. Exact (min weight) line shows the calculation using the repetition code based formulas in the text, without any approximations. For the numerical data, each data point is based on $10^{4}$ samples.
FIGURE 8: Logical infidelity $r_{1}$ for $Z$-Shor codes in a low error regime. It is clear that there is no true threshold, with the crossing between $d,d+2$ curves decreasing with increasing distance. Exact (min weight) line shows the calculation using the repetition code based formulas in the text, without any approximations. For the numerical data, each data point is based on $10^{4}$ samples. The low amount of samples are responsible for the large deviations at low error rates from the analytical formulae.
FIGURE 9: Logical infidelity $r_{1}$ for $Z$-Shor codes in a low error regime. It is clear that there is no true threshold, with the crossing between $d,d+2$ curves decreasing with increasing distance. Exact (min weight) line shows the calculation using the repetition code based formulas in the text, without any approximations. For the numerical data, each data point is based on $10^{4}$ samples. The low amount of samples are responsible for the large deviations at low error rates from the analytical formulae.
FIGURE 10: Logical infidelity $r_{1}$ for $C_{l,2}$ Z-stacked Shor codes. Exact (min weight) line shows the calculation using the repetition code based formulas in the text, without any approximations. For the numerical data, each data point is based on $10^{4}$ samples. The low amount of samples are responsible for the large deviations at low error rates from the analytical formulae.
FIGURE 11: Logical infidelity $r_{1}$ for $C_{l,3}$ Z-stacked Shor codes. Exact (min weight) line shows the calculation using the repetition code based formulas in the text, without any approximations. For the numerical data, each data point is based on $10^{4}$ samples. The low amount of samples are responsible for the large deviations at low error rates from the analytical formulae.